# Operation of a dual-phase xenon detector with wavelength sensitivity from ultraviolet to infrared

### v1

@@ 3 Introduction
Around the year 2000, the first studies of xenon’s IR continuum emission identified a broad scintillation peak centered around 1.3 m (Borghesani et al. (2001)). Initial measurements were conducted in gaseous xenon (Carugno (1998)), where the IR light yield was found to be comparable in magnitude to that of the UV scintillation (Belogurov et al. (2000)). A blue shift of the emission peak with increasing pressure was also observed and later modeled (Borghesani et al. (2007a, b, 2025)). Above an electric field of about 500 V/cm in the gas, the IR signal was found to increase roughly in proportion to the ionization charge, consistent with a signal from electroluminescence (Belogurov et al. (2000)). IR scintillation has additionally been detected in liquid xenon (Bressi et al. (2000)), though with a light yield estimated to be two orders of magnitude lower than in gas. Furthermore, there are indications that the emission spectrum in the liquid is significantly shifted toward shorter wavelengths (Bressi et al. (2001)).
To better understand xenon’s IR scintillation properties, we initiated a dedicated R&D program. As a first step, we operated a room-temperature setup using gaseous xenon near atmospheric pressure. Using this setup, we studied the time response and light yield of IR scintillation following -particle irradiation, under varying pressure and gas purity. We observed that at approximately 1 bar, the light yield is about one-third of that in the UV. We found that the time response is well described by a sum of two exponential decays: a fast decay with a time constant of , contributing less than 1 % of the total yield, and a slower decay with a time constant of (Piotter et al. (2023); Hammann et al. (2024)). We observed a linear increase in light yield with pressure between 500 and 1100 mbar, along with a decrease in yield as purity decreased. These measurements were performed using a PMT sensitive to wavelengths between about 950 and 1700 nm.
@@ 4.3 The HeXe gas handling system
The xenon purity is assessed by measuring the “electron lifetime”, a parameter that quantifies the loss of drifting electrons in the liquid due to attachment to electronegative impurities. During this measurement campaign, the electron lifetime varied from about 40 to 50 s, notably shorter than in previous runs (a few hundred to a few thousand s). This reduction was likely caused by outgassing from the IR PMT materials and the PTFE structure of the TPC.
5 Calibration source, DAQ, and data analysis
@@ 6.1 Signal time structure
The fastest decay component is not well resolved due to the time smearing and is therefore simply modeled by a Gaussian. Its timescale may be comparable to the few-ns component previously observed in gaseous xenon at room temperature (Piotter et al. (2023); Hammann et al. (2024)). The second fastest decay time of approx. s is also similar to the decay time found in previous measurements. The third component of s decay time has not been reported before. This slowest component contributes more than 50% of the total signal integral, making it the dominant contribution. Interestingly, this component is very similar to the slow component found after the (purple line in fig. 5), pointing to a common origin. The fastest component accounts for approximately one-third of the total signal.
@@ 6.2 Electric field dependence
To study the IR emission following for different electron energies, we varied the amplification field in the xenon gas region from about 9 kV/cm to 15 kV/cm by varying the anode potential, while keeping the drift field constant. The IR PMT bias voltage was constant across all measurements.
In order to estimate the liquid level in the TPC, signals from the photoemission of the gate electrode for different amplification fields are employed. The liquid level is then matched using the gate-surface times difference and the field-dependent drift velocity in the liquid (from Gushchin et al. (1982)). Averaging over all anode voltages and two analysis methods yields a liquid level of mm above the gate. The uncertainty is propagated to the electric fields in the liquid, which directly affect the extraction efficiency. Table 1 summarizes the applied voltage differences, along with the calculated electric fields in the liquid and gas phases, based on the derived liquid level.
@@
The amplification fields studied here cover a region where the dependence of the UV signal strength on the field strength transitions from a linear to an exponential dependence (Aprile et al. (2014)). We observe a monotonic increase with the field strength in the IR photon rate, as shown in fig. 8 (top panel). A similar trend was also observed in Belogurov et al. (2000) for lower fields.
@@ 7 Discussion and outlook
The IR-sensitive PMT used for the measurements presented here is also sensitive to shorter wavelengths as shown in fig. 2, adding some uncertainty to the interpretation of the results. In particular, it is sensitive to the tail of the third continuum emission (see fig. 2), which has a decay constant of approximately 8 ns (Millet et al. (1978)). This component has so far been measured as part of the primary scintillation light in gaseous xenon for -particles at pressures up to 10 bar, as well as for electrons at 10 bar (Leardini et al. (2022)). In electroluminescence, it was shown that the third continuum contributes at most at the sub-percent level to . Given the PMT’s roughly threefold lower QE at this wavelength and partial spectral overlap, its contribution is considered subdominant. An additional contribution spanning a broad wavelength range arises from neutral bremsstrahlung (Henriques et al. (2022); Leardini et al. (2022)). To the best of our knowledge, neither of these processes has been systematically studied in the liquid phase.
@@
In the IR time response following both and signals, we observed a time constant of approximately 30 s, which accounts for roughly half of the respective total signals. This similarity may point to a common origin of the slow component. Delayed electron tails have been observed in several liquid xenon TPCs (see for example Aprile et al. (2014)). A fraction of liberated electrons are not immediately extracted into the gas phase. Instead, they are reflected off the liquid surface and eventually extracted at later times. In addition, – in particular when the xenon purity is not particularly high – UV scintillation light may photoionize residual impurities in the liquid, generating delayed electrons. However, any signal from delayed electrons would also be visible in the UV channel, and thus can only account for the tail observed in the signal. Another possible contribution may arise from material fluorescence. The TPC wall is made of PTFE, which has been suggested as a source of fluorescent light (Shaw et al. (2007))), although it is debated (Araujo et al. (2019)) or attributed instead to fluorescence of residues in the PTFE (Pollmann (2025)). Fluorescence from quartz – the material used for the UV PMT window – has also been reported to produce a long signal tail in the UV-sensitive detectors (Sorensen and Gibbons (2025)), but could also contribute to the slow component observed in the IR channel. We also cannot rule out a contribution from a PMT-related effect, possibly induced by the suboptimal bias voltage applied to the photocathode, which may lead to delayed photoelectron extraction from the InP/InGaAs photocathode.
@@
Although our data is not sensitive to the time constant of the fast IR scintillation component in the gas, it is sensitive to the amplitude of this component. Our past measurements at room temperature showed that, for -particles, the ns component contributes less than 1% to the total signal (Hammann et al. (2024)). The current measurements with electrons in the gas show that the fast component makes up about one-third of the signal. This difference could potentially allow particles to be discriminated in the gas by means of their pulse shape.

### v2

@@ 3 Introduction
A first systematic investigation of IR scintillation in noble gases was conducted by Lindblom and Solin (1988). Around the year 2000, the first studies of xenon’s IR continuum emission identified a broad scintillation peak centered around 1.3 m (Borghesani et al. (2001)), consistent with earlier predictions from potential curve estimates (Mulliken (1970)). Initial measurements were conducted in gaseous xenon (Carugno (1998)), where the IR light yield was found to be comparable in magnitude to that of the UV scintillation (Belogurov et al. (2000)). A red shift of the emission peak with increasing pressure was also observed and later modeled (Borghesani et al. (2007a, b, 2025)). Above an electric field of about 500 V/cm in the gas, the IR signal was found to increase roughly in proportion to the ionization charge, consistent with a signal from electroluminescence (Belogurov et al. (2000)). IR scintillation has additionally been detected in liquid xenon (Bressi et al. (2000)), though with a light yield estimated to be two orders of magnitude lower than in gas. Furthermore, there are indications that the emission spectrum in the liquid is significantly shifted toward shorter wavelengths (Bressi et al. (2001)).
To better understand xenon’s IR scintillation properties, we initiated a dedicated R&D program. As a first step, we operated a room-temperature setup using gaseous xenon near atmospheric pressure. Using this setup, we studied the time response and light yield of IR scintillation following -particle irradiation, under varying pressure and gas purity. We observed that, at approximately 1 bar, the light yield is about one-quarter of that in the UV. We found that the time response is well described by a sum of two exponential decays: a fast decay with a time constant of , contributing less than 1 % of the total yield, and a slower decay with a time constant of (Piotter et al. (2023); Hammann et al. (2024)). We observed a linear increase in light yield with pressure between 500 and 1100 mbar, along with a decrease in yield as purity decreased. These measurements were performed using a PMT sensitive to wavelengths between about 950 and 1700 nm.
@@ 4.3 The HeXe gas handling system
The xenon purity is assessed by measuring the “electron lifetime”, a parameter that quantifies the loss of drifting electrons in the liquid due to attachment to electronegative impurities. During this measurement campaign, the electron lifetime varied from about 40 to 50 s. This corresponds to an -equivalent mole fraction of approximately using the electron attachment rate constant from Bakale et al. (1976). The electron lifetime is notably shorter than in previous runs (a few hundred to a few thousand s), which was likely caused by outgassing from the IR PMT materials and the PTFE structure of the TPC.
5 Calibration source, DAQ, and data selection
@@ 6.1 Signal time structure
The fastest decay component is not well resolved due to the time smearing and is therefore simply modeled by a Gaussian. Its timescale may be comparable to the few-ns component previously observed in gaseous xenon at room temperature (Piotter et al. (2023); Hammann et al. (2024)). The second fastest decay time of approx. s is also similar to the decay time found in previous measurements. The third component of s decay time has not been reported in xenon before. This slowest component contributes more than 50% of the total signal integral, making it the dominant contribution. Interestingly, this component is very similar to the slow component found after the (purple line in fig. 5), pointing to a common origin. The fastest component accounts for approximately one-third of the total signal.
@@ 6.2 Electric field dependence
To study the IR emission following for different electron energies, we varied the amplification field in the xenon gas region from about 9 kV/cm to 15 kV/cm by varying the anode potential and keeping the drift field constant. The IR PMT bias voltage was constant across all measurements.
An iterative procedure was employed to determine the liquid level. Electrons generated by photoionization on the gate via bright UV signals were used, which have a drift distance corresponding to the liquid level. Since the electron drift velocity depends on the electric field (Gushchin et al. (1982)), which itself is a function of the liquid level, the assumed level was adjusted iteratively until the calculated and observed drift times between the gate and the surface matched. Averaging over all anode voltages and two analysis methods yielded a liquid level of mm above the gate. The uncertainty is propagated to the electric fields in the liquid, which directly affect the extraction efficiency. Table 1 summarizes the applied voltage differences, along with the calculated electric fields in the liquid and gas phases, based on the derived liquid level.
@@
We observe a monotonic increase with the field strength in the IR photon rate, as shown in fig. 8 (top panel). A similar trend was also observed in Belogurov et al. (2000) for lower fields.
@@ 7 Discussion and outlook
The IR-sensitive PMT used for the measurements presented here is also sensitive to shorter wavelengths as shown in fig. 2, adding some uncertainty to the interpretation of the results. In particular, it is sensitive to the tail of the third continuum emission, which has a decay constant of approximately 8 ns (Millet et al. (1978)). This component has so far been measured as part of the primary scintillation light in gaseous xenon for -particles at pressures up to 10 bar, as well as for electrons at 10 bar (Leardini et al. (2022)). In electroluminescence, it was shown that the third continuum contributes at most at the sub-percent level to . Given the PMT’s roughly threefold lower QE at this wavelength and partial spectral overlap, its contribution is considered subdominant. An additional contribution spanning a broad wavelength range arises from neutral bremsstrahlung (Henriques et al. (2022); Leardini et al. (2022)). To the best of our knowledge, neither of these processes has been systematically studied in the liquid phase.
@@
In the IR time response following both and signals, we observed a time constant of approximately 30 s, which accounts for roughly half of the respective total signals. This similarity may point to a common origin of the slow component. Notably, a comparable slow component has been observed in the primary IR scintillation of both liquid and gaseous argon (Buzulutskov et al. (2011); Bondar et al. (2012)) with unresolved origin, though it may be attributed to a slow scintillation process. Nevertheless, alternative explanations should also be considered. Delayed electron tails have been observed in several liquid xenon TPCs (see for example Aprile et al. (2014)). A fraction of liberated electrons are not immediately extracted into the gas phase. Instead, they are reflected off the liquid surface and eventually extracted at later times. In addition – in particular when the xenon purity is not particularly high – UV scintillation light may photoionize residual impurities in the liquid, generating delayed electrons. However, any signal from delayed electrons would also be visible in the UV channel, and thus can only account for the tail observed in the signal. Another possible contribution may arise from material fluorescence. The TPC wall is made of PTFE, which has been suggested as a source of fluorescent light (Shaw et al. (2007))), although it is debated (Araujo et al. (2019)) or attributed instead to fluorescence of residues in the PTFE (Pollmann (2025)). Fluorescence from quartz – the material used for the UV PMT window – has also been reported to produce a long signal tail in the UV-sensitive detectors (Sorensen and Gibbons (2025)), but could also contribute to the slow component observed in the IR channel. We also cannot rule out a contribution from a PMT-related effect, possibly induced by the suboptimal bias voltage applied to the photocathode, which may lead to delayed photoelectron extraction from the InP/InGaAs photocathode.
@@
Although our data is not sensitive to the time constant of the fast IR scintillation component in the gas, it is sensitive to the amplitude of this component. Our past measurements at room temperature showed that, for -particles, the ns component contributes less than 1% to the total signal (Hammann et al. (2024)). The current measurements with electrons in the gas show that the fast component makes up about one-third of the signal. This difference may result from variations in the population of excited states caused by differing ionization densities. As a result, IR pulse shape discrimination could be a viable method for identifying particle types in the gas.